The largest wavelength in the uv region of the hydrogen spectrum is 122 nm. the smallest wavelength in the infrared region of the hydrogen spectrum is?
a). 802nm
b). 823nm
c). 1882nm
d). 1648nm

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
(B) option seemss corect.
1/l=R(1/n22-1/n­12)
=>1/122=R(1-1/4)
=>1/122=3R/4
=>R=4/3x122

smallest l in the infrared region. n1= infinity
n2=3
Therefore, 1/l = R(1/32-1/infinity)
=R/9
l=9/R=823nm.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)